zc01638:

Anastasia uses the equation p = 0.7(rh + b) to estimate the amount of take-home pay, p, for h hours worked at a rate of r dollars per hour and any bonus received, b. What is an equivalent equation solved for h?

Vocaloid:

p = 0.7(rh + b) We want to isolate h so start by dividing both sides by 0.7 p/0.7 = rh + b Subtract b then divide both sides by r